How will you compare the range (horizontal distance) travelled by the

ball at angles of 25º and 65º? How about at angles of 40º and 50º?


1
Expert's answer
2021-06-02T09:37:16-0400

We know that according to the equation


"R=\frac{v^2\sin(2\theta)}{g}"

the maximum range is at 45°. Compare the ranges at 25° and 65°:


"R(25°)=\frac{v^2\sin(2·25°)}{g}=0.766\frac{v^2}{g},\\\space\\
R(65°)=\frac{v^2\sin(2·65°)}{g}=0.766\frac{v^2}{g},\\\space\\
R(25°)=R(65°)."

For 40° and 50°:


"R(40°)=\frac{v^2\sin(2·40°)}{g}=0.985\frac{v^2}{g},\\\space\\
R(50°)=\frac{v^2\sin(2·50°)}{g}=0.985\frac{v^2}{g},\\\space\\
R(40°)=R(50°)."